How do I figure out the standard form of a quadratic equation?
y = -X^2 +2X + 13
I got -(X-2)^2 + 15 ....but that is wrong..?
Help me find out what I am doing wrong? Thanks!

Answer by stanbon(75887)

y = -X^2 +2X + 13
Note:: That is already in standard form.
-----
Maybe you want it in intercept form::
x^2 -2x = 13-y
-----
x^2 -2x + 1 = 14-y
-----
(x+1)^2 = -(y-14)
===================
Cheers,
Stan H.